OpenStudy (anonymous):

the number of grains of sand in one cubic centermeter sample was 9.37 x 10 cubed. how many grains of sand would be in a cubic meter of sand

OpenStudy (whpalmer4):

This is a basic unit conversion problem. You need to know that 1 m = 100 cm, or expressed as a fraction, \[\frac{1\text{ m}}{100\text{ cm}} =1\] Now put that in the expression: \[9.37*10^3 \frac{\text{grains}}{ \text{ cm}^{3}} * 1 \text{m}^3 * (\frac{100\text{ cm}}{1\text{ m}})^3\]Notice how the units of length cancel out, leaving only grains.

OpenStudy (anonymous):

so there's no answer or is there an answer

OpenStudy (whpalmer4):

Evaluate that expression... 9.37 * 10^3 * 1 * (100/1)^3 = ?

OpenStudy (whpalmer4):

\[9.37*10^3 \frac{\text{grains}}{ \text{ cm}^{3}} * 1 \text{m}^3 * (\frac{100\text{ cm}}{1\text{ m}})^3= \]\[9.37*10^3\frac{\text{grains}}{ \text{ cm}^{3}}*1\text{m}^3 *100^3\text{ cm}^3/(1^3 \text{ cm}^3)=\]\[9.37*10^9 \text { grains} \]

OpenStudy (anonymous):

ohhh thanks(: